# Jingjing Sun, Pengyi Shi, Jim Dai

### v3

@@
Problem Definition: Managing inpatient flow in large hospital systems is challenging due to the complexity of assigning randomly arriving patients – either waiting for primary units or being overflowed to alternative units. Current practices rely on ad-hoc rules, while prior analytical approaches struggle with the intractably large state and action spaces inherent in patient-unit matching. A scalable decision-support framework is needed to optimize overflow management while accounting for time-periodic fluctuations in patient flow. Methodology/Results: We develop a scalable decision-making framework using Proximal Policy Optimization (PPO) to optimize overflow decisions in a time-periodic, long-run average cost setting. To address the combinatorial complexity, we introduce atomic actions, which decompose multi-patient routing into sequential, tractable assignments. We further enhance computational efficiency through a partially-shared policy network designed to balance parameter sharing with time-specific policy adaptations, and a queueing-informed value function approximation to improve policy evaluation. Our method significantly reduces the need for extensive simulation data, a common limitation in reinforcement learning applications. Case studies on hospital systems with up to twenty patient classes and twenty wards demonstrate that our approach matches or outperforms existing benchmarks, including approximate dynamic programming, which is computationally infeasible beyond five wards. Managerial Implications: Our framework offers a scalable, efficient, and adaptive solution for managing patient flow in complex hospital systems. More broadly, our results highlight that domain-aware adaptation – leveraging queueing structures and operational insights – is more critical to improving algorithm performance than fine-tuning neural network parameters when applying general-purpose algorithms to specific domain applications.
@@
[…]
@@ 1.1 Paper Overview and Contribution
This paper addresses the existing gap by developing an algorithmic framework based on the Proximal Policy Optimization (PPO) method to minimize the long-run average costs associated with overflow. Our key contributions are summarized as follows.
@@
    The success of this framework hinges on two key elements: (a) leveraging the “atomizable” action space and (b) incorporating domain knowledge of queueing systems – particularly in designing value function approximation and policy representation. We elaborate on them next.
@@
    Action Atomization: We propose the atomic action in Section 4.1, which decomposes class-level decisions into sequential individual-level assignments. This transformation fundamentally simplifies the original combinatorial action space and allows for efficient policy optimization via gradient-based methods. Despite its seemingly myopic nature, we demonstrate in our case study that this approach achieves performance comparable to brute-force optimization over the entire action space (Dai and Shi 2019), i.e., this approach works. Moreover, it retains flexibility for aggregate planning and incorporating individual preferences for practical implementation.
@@
    Dealing with Time-periodicity: To account for real-world daily and weekly periodic patterns in patient arrivals and discharges, we integrate time-periodicity into our PPO framework by leveraging the queueing system structure. First, we establish a theoretical guarantee for policy improvement of PPO under the periodic setting (Section 4.2). Second, we design a partially-shared policy network that balances parameter sharing and temporal specificity (Section 5.1). This design captures structural similarities while being adaptive to time variations. It outperforms two conventional alternatives: (i) a fully-shared network that requires more complex policy networks to model time dependencies, and (ii) a fully-separate network that demands significantly more training data. Lastly, we propose a new value function approximation derived from the Poisson equation and tailored to the queueing dynamics under the atomic actions – it decomposes the system-wide value function into pool-wise approximations to improve computational efficiency and stability (Section 5.2). Case study results in Section 6 confirm that our framework, which integrates all these design elements, enhances sample efficiency while maintaining superior performance over other benchmarks.
@@ 2.1 Bed Partition and Patient Routing
Adding new beds to alleviate inpatient congestion is often not feasible due to constraints related to space, finances, and regulations. Therefore, we focus on literature that explores strategies for utilizing existing bed resources more efficiently. Best et al. (2015) examine bed management at a strategic level, where the decision involves partitioning inpatient beds into separate wings dedicated to different specialties (overflow is not allowed). As a follow-up, Izady and Mohamed (2021) consider a similar setting but introduce the option of “pooled wings,” which can accommodate patients from multiple specialties when dedicated wings reach capacity. Optimization techniques are primarily employed in this type of strategic planning to maximize inpatient bed utilization.
@@ 2.2 PPO Applications
In conventional PPO algorithms, the policy space is directly parameterized by an NN, with the output size corresponding to the number of possible actions – typically small in other applications (e.g., Schulman et al. (2017) and Dai and Gluzman (2022) focus on smaller action spaces). However, in our problem, the action space is significantly larger, making it computationally prohibitive to train the policy network using traditional parameterization methods. To our knowledge, there are no established examples in the literature where PPO has been successfully applied to such large action spaces, except for Feng et al. (2021), which showcase an effective use of action decomposition in ride-hailing applications. However, their setting is finite-horizon, whereas the framework we develop in this paper tackles the more complex infinite-horizon problem. Additionally, to account for the periodic nature of our infinite-horizon MDP, we introduce a novel partially-shared network structure that outputs actions at different epochs throughout the day. This design is conceptually related to specialized architectures in federated learning (Arivazhagan et al. 2019, Kulkarni et al. 2020), which address statistical heterogeneity, a different problem from ours.
@@ 3 Model Formulation
In this section, we present the queueing model formulation abstracted from the overflow problem. While we describe a general queueing model, we ensure it retains the key features relevant to our application. Specifically, the model is based on a parallel server system with multiple customer classes and server pools. Each customer class represents a medical specialty (or a cluster of related specialties), and each server pool corresponds to an inpatient unit (ward), with servers representing individual inpatient beds. We detail the queueing discipline along with the arrival and departure processes in Section 3.1. We formulate the overflow decision problem as an MDP in Section 3.2. Table 5 in Appendix 8 summarizes all the notations introduced in the model formulations and those used in later sections.
@@ 3.1 Parallel-server System
[…]
@@ 3.2 MDP Model
We use an MDP framework to decide on whether to assign a waiting customer to one of the overflow pools. The objective is to find the optimal overflow policy that minimizes the infinite-horizon, long-run average cost. We consider a discrete-time setup, where the decision maker (e.g., hospital manager) makes overflow decisions at $m$ predefined time epochs each day (e.g., hourly), denoted as $t_{0},t_{1},\dots,t_{m-1}$ for day $t$. In the rest of this section, we define the state, action, cost, policy, and objective of the MDP.
@@
  • $X_{j}(t_{k})$, $j=1,\dots,J$, denotes the customer count for pool $j$. This customer count includes both the number of customers waiting in buffer $j$ and those in service in pool $j$; we show below on how to recover each part. The former part (waiting count) only includes class $j$ customers, while the latter (in-service count) may include customers from other classes due to overflow assignments.
  • $Y_{j}(t_{k})$, $j=1,\dots,J$, denotes the to-depart count in pool $j$. This quantity represents the number of customers expected to depart from pool $j$ between the present time and the end of the current day. Tracking this quantity is essential for the two-time-scale inpatient departure scheme, where the “planned discharge” information is used to inform overflow decisions.
@@
To explain, the total number of overflow assignments from all non-primary classes to pool $j$ must not exceed the number of idle servers in pool $j$, which can be computed as $(N_{j}-x_{j})\vee 0$, with $x\vee 0=\max(x,0)$ for any real number $x$. Similarly, the total number of assignments for class $i$, including overflow assignments to a pool $\ell\neq i$ and those kept waiting, must equal the number of waiting customers. This can be calculated as $(x_{i}-N_{i})\vee 0$, which we also refer to as the queue length. We use $q_{i}=(x_{i}-N_{i})\vee 0$ to denote this queue length for class $i$ (buffer $i$).
@@
We denote the state at epoch $t_{k}$ before taking an action as $s$, i.e., the pre-action state. After an action is taken, the resulting state is denoted as $s^{+}$, i.e., the post-action state. Given an action $f={f_{i,j}}$, the transition from $s$ to $s^{+}$ is as follows:
$x^{+}_{j}$ $=x_{j}+\sum_{i=1,i\neq j}^{J}f_{i,j}-\sum_{\ell=1,\ell\neq j}^{J}f_{j,\ell},\quad j=1,\dots,J;$ (2)
$y_{j}^{+}$ $=y_{j},\quad j=1,\dots,J;$
$h^{+}$ $=h.$
We denote the state, after the exogenous events (random arrivals and departures) occurring between the current epoch $t_{k}$ and the next epoch $t_{k+1}$, as $s^{\prime}$. The corresponding transition probability from state $s$ to state $s^{\prime}$, through $s^{+}$ under action $f$, is denoted as $p(s^{\prime}|s,f)$. The detailed calculation of these transition probabilities is provided in Appendix 9.
@@
where $q_{j}^{+}=(x_{j}^{+}-N_{j})\vee 0$ denotes the post-action queue length for class $j$, and the transition to the post-action state follows (2). This cost setup is designed to identify a policy that strikes a balance between the negative effects of waiting and the undesirability of overflow assignments.
@@
We consider the long-run average cost objective since overflow decisions have long-term impacts on the system. We focus on all feasible stationary Markovian policies, and we denote the policy space as $\Pi$. For a policy $\pi\in\Pi$, we use $\pi(f|s)$ to denote the probability of selecting a feasible action $f$ in a given state $s$ (where deterministic policies are included via using the Dirac measure). The resulting stochastic process $\{S(t_{k})\}$ forms a discrete-time Markov Chain (DTMC) with a one-epoch transition probability from state $s$ to $s^{\prime}$ as $p_{\pi}(s^{\prime}|s)=\mathbb{E}_{f\sim\pi(\cdot|s)}p(s^{\prime}|s,f)$. The expected one-epoch cost is $g_{\pi}(s)=\mathbb{E}_{f\sim\pi(\cdot|s)}g(s,f)$. Correspondingly, the objective is to identify the optimal policy
@@
While our baseline setting does not consider transferring overflow customers back to their primary pool, this can be accommodated by adding actions that represent the number of in-service transfers from each pool $j$ back to their primary pools. In addition, although we assume a linear cost function in this paper, our framework can handle more complex cost structures, such as polynomial costs or costs related to service levels (e.g., the proportion of patients waiting more than 6 hours). See the appendix of Dai and Shi (2019) for an example of modifying the cost to account for service-level metrics. We note that these adjustments do not alter the core elements of our problem, and our algorithmic framework is both flexible and general enough to accommodate them. For ease of exposition, throughout the remainder of the paper, we will focus on the current baseline setting, where primary assignments are handled automatically, departures are pool-dependent, no transfer back of overflow customers, and the cost is linear.
@@ 4 Atomic Actions and Proximal Policy Optimization
One of the core challenges in solving the MDP model developed in Section 3.2 is the combinatorial explosion of the action space. When assigning servers for multiple customers simultaneously, the number of possible combinations grows exponentially with the number of customer classes and server pools. Specifically, when there are enough idle servers in the system, the action space has a size of $\prod_{j=1}^{J}\left(\begin{array}[]{c}q_{j}+w_{j}\\ w_{j}\end{array}\right)$, where $q_{j}$ is the number of waiting customers in class $j$ and $w_{j}$ is the number of overflow pools for class $j$ that have idle servers. To illustrate, in a twenty-pool system with each server-pool having a capacity of 60 and a state $(x_{1},x_{2},x_{3},\dots,x_{20})=(65,63,62,50,50,50,\dots,50)$, the number of potential ways to assign 10 waiting customers can exceed $10^{9}$. In other words, even if the value function is computed (either exactly or via approximation), brute-force searching through all possible actions to find the one with the lowest cost-to-go is infeasible. To overcome this curse of dimensionality, we use atomic actions in combination with PPO to directly optimize the policy, which forms the key foundation for our algorithm’s scalability. In Section 4.1, we introduce the atomic action and policy setup. In Section 4.2, we present the framework of the PPO algorithm.
@@ 4.1 Action Space Reduction: Atomic Actions
To elaborate, for a given state $s=(x_{1},\dots,x_{J},y_{1},\dots,y_{J},h)$ observed at epoch $t_{k}$, the total number of waiting customers across all classes is $q=\sum_{j=1}^{J}q_{j}$, where $q_{j}=(x_{j}-N_{j})\vee 0$. Following a pre-determined rule (e.g., first-come-first-serve), we determine an order for these $q$ waiting customers and conduct atomic actions $\mathbf{a}=\{a^{1},a^{2},\dots,a^{q}\}$ for each of them sequentially. Here, $a^{n}$ represents the assignment for the $n$th customer in the chosen order. The action space for each $a^{n}$ is simply $\mathcal{J}=\{1,\dots,J\}$. If the class of the $n$th customer is $j$, then $a^{n}=j$ indicates the customer will continue waiting until the next decision epoch; otherwise, the customer is assigned to an overflow pool (see more details on feasible overflow assignment in Section 4.1.2).
$s^{1}=s$$s^{\prime}$$\textbf{a}=(a^{1},\dots,a^{q})$$s^{2}$$s^{3}$$s^{q+1}$$\sigma^{1}(s),a^{1}$$\sigma^{2}(s),a^{2}$rrivalDepartureTime $t_{k}$ime $t_{k+1}$
@@
Figure 1 illustrates the atomic decision process for deciding and executing the atomic actions. We denote the sequence of customer classes of the chosen order as $\boldsymbol{\sigma}(s)=(\sigma^{1}(s),\sigma^{2}(s),\cdots,\sigma^{q}(s))$, with $\sigma^{n}(s)$ representing the class of the $n$th waiting customer within the chosen order. Starting from the first customer, who is from class $\sigma^{1}(s)$, we execute the atomic action $a^{1}$ for this customer, then we update the state according to $a^{1}$ and sequentially move to the second customer, so on and forth till we execute the atomic action $a^{q}$ for the last waiting customer, who is from class $\sigma^{q}(s)$. The atomic action $a^{n}$ for the $n$th customer is chosen based on the atomic state, $s^{n}$, which is observed after assigning the $(n-1)$th customer. This state has the same structure as the one introduced in Section 3.2, i.e., $s^{n}=(x^{n}_{1},\dots,x^{n}_{J},y_{1},\dots,y_{J},h)$. After executing $a^{n}$, the atomic decision process updates the state to $s^{n+1}$ as
@@
For clarity, and to distinguish between the MDP state and action defined in Section 3.2 and those used for the atomic decision process here, we will refer to the state and action in Section 3.2 as the system-level state and action throughout the rest of the paper. Meanwhile, we will refer to the state and action in the atomic decision process as the atomic state and action. It is important to note that the system-level and atomic states and actions are closely related. Specifically, the initial atomic state $s^{1}$ corresponds to the system-level (pre-action) state $s$, and after all atomic actions are executed, the final atomic state $s^{q+1}$ corresponds to the system-level post-action state $s^{+}$. Additionally, the sum of the atomic actions for the inflow and outflow of class $j$ customers can be used to recover the system-level actions. That is, for any given pair $(\boldsymbol{\sigma}(s),\textbf{a})$, we can recover the system-level action $f_{i,j}$ via $\sum_{n=1}^{q}\mathbbm{1}(a^{n}=j,\sigma^{n}(s)=i)$.

### v4

@@
Problem Definition: Managing inpatient flow in large hospital systems is challenging due to the complexity of assigning randomly arriving patients – either waiting for primary units or being overflowed to alternative units. Current practices rely on ad-hoc rules, while prior analytical approaches struggle with the intractably large state and action spaces inherent in patient-unit matching. Scalable decision support is needed to optimize overflow management while accounting for time-periodic fluctuations in patient flow. Methodology/Results: We develop a scalable decision-making framework using Proximal Policy Optimization (PPO) to optimize overflow decisions in a time-periodic, long-run average cost setting. To address the combinatorial complexity, we introduce atomic actions, which decompose multi-patient routing into sequential assignments. We further enhance computational efficiency through a partially-shared policy network designed to balance parameter sharing with time-specific policy adaptations, and a queueing-informed value function approximation to improve policy evaluation. Our method significantly reduces the need for extensive simulation data, a common limitation in reinforcement learning applications. Case studies on hospital systems with up to twenty patient classes and twenty wards demonstrate that our approach matches or outperforms existing benchmarks, including approximate dynamic programming, which is computationally infeasible beyond five wards. Managerial Implications: Our framework offers a scalable, efficient, and explainable solution for managing patient flow in complex hospital systems. More broadly, our results highlight that domain-aware adaptation is more critical to improving algorithm performance than fine-tuning neural network parameters when applying general-purpose algorithms to specific applications.
@@
Time-Periodic Multi-class Queue; Patient Flow; Scalable Decision-Making
@@ 1.1 Paper Overview and Contribution
This paper addresses the existing gap by developing an algorithmic framework based on the Proximal Policy Optimization (PPO) method. Our key contributions are summarized as follows.
@@
    The success of this framework hinges on two key elements: (a) leveraging the “atomizable” action space and (b) incorporating domain knowledge of queueing systems.
@@
    Action Atomization: As detailed in Section 4.1, atomic action decomposes class-level decisions into sequential individual-level assignments, which is critical for simplifying the original combinatorial action space. Despite its seemingly myopic nature, we show that this approach, when combined with PPO, achieves performance comparable to the ADP benchmark in a five-pool system, the largest setting considered in Dai and Shi (2019); i.e., this approach works. The atomic formulation is inspired by Feng et al. (2021), who coined the term atomic action decomposition in the context of finite-horizon ride-hailing problems. Extending this idea to the more challenging infinite-horizon, long-run average cost setting is nontrivial, and to the best of our knowledge, this paper is the first to show that combining atomic actions with PPO is effective in this regime. We discuss related developments in this growing research area in Section 2.
@@
    Dealing with Time-periodicity: To account for real-world daily and weekly periodic patterns in patient arrivals and discharges, we integrate time-periodicity into our PPO framework. First, we establish a theoretical guarantee for policy improvement of PPO under the periodic setting (Section 4.2). Second, we design a partially-shared policy network that balances parameter sharing and temporal specificity (Section 5.1). Lastly, we propose a new value function approximation tailored to the queueing dynamics under the atomic actions to improve computational efficiency (Section 5.2).
  4. 4.
    Actionable Insights: Case study results in Section 6, including ablation studies, confirm that our integrated framework enhances sample efficiency while maintaining superior performance over other benchmarks. These results highlight that domain-aware algorithm design is more impactful than tuning generic hyperparameters. Moreover, our policy visualization and interpretability analysis reveal that the learned policies are both intuitive (aligning with managerial expectations) and insightful (uncovering nuanced network effects that are difficult to capture with heuristic rules). These insights provide actionable guidelines for managing overflow while also highlight the necessity of sophisticated, data-driven tools in complex hospital networks.
@@ 2.1 Bed Partition and Patient Routing
Adding new beds to alleviate inpatient congestion is often not feasible due to constraints related to space, finances, and regulations. Therefore, we focus on literature that explores strategies for utilizing existing bed resources more efficiently (Zamani et al. 2024). Best et al. (2015) examine bed management at a strategic level, where the decision involves partitioning inpatient beds into separate wings dedicated to different specialties (overflow is not allowed). As a follow-up, Izady and Mohamed (2021) consider a similar setting but introduce the option of “pooled wings,” which can accommodate patients from multiple specialties when dedicated wings reach capacity.
@@ 2.2 PPO Applications
In conventional PPO algorithms, the policy space is directly parameterized by a Neural Network (NN), with the output size corresponding to the number of possible actions, typically a small number (e.g., Schulman et al. (2017), Dai and Gluzman (2022)). However, in our problem, the action space is significantly larger, making it computationally prohibitive to train the policy network using traditional parameterization methods. Feng et al. (2021) demonstrates the effectiveness of combining action decomposition with PPO for ride-hailing, but only in a finite-horizon setting. While PPO has been successfully applied to infinite-horizon MDPs (Dai and Gluzman 2022), to our knowledge, there are no established examples of its application to problems combining both large action spaces and infinite-horizon objectives. Our work addresses both challenges simultaneously, handling large action spaces in a time-varying, infinite-horizon environment. The study of atomic-action-based PPO is actively growing. While this paper was under review, concurrent work by Huo (2025) and Dai et al. (2025a) have demonstrated the effectiveness of this framework in data switch scheduling, and in battery charging and electric vehicle scheduling for robo-taxi systems, respectively; the latter introduced the term Atomic-PPO. Motivated by the empirical success, Dai et al. (2025b) further provides a theoretical justification for atomic action decomposition.
@@ 3 Model Formulation
In this section, we present the queueing model formulation abstracted from the overflow problem. While we describe a general queueing model, we ensure it retains the key features relevant to our application. Specifically, the model is based on a parallel server system with multiple customer classes and server pools. Each customer class represents a medical specialty (or a cluster of related specialties), and each server pool corresponds to an inpatient unit (ward), with servers representing individual inpatient beds. We detail the queueing discipline along with the arrival and departure processes in Section 3.1. We formulate the overflow decision problem as an MDP in Section 3.2. Appendix 8 summarizes all the notations.
@@ 3.1 Parallel-server System
Arrival process. The underlying system evolves on a continuous-time scale. We assume that the arrival process for each customer class follows a periodic Poisson process, with the associated arrival rate function $\lambda_{j}(t)$ being periodic with a period of $T$ for each $j\in\mathcal{J}$. For ease of exposition, we set $T=1$ day to capture the time-varying pattern within a day. The corresponding daily arrival rate for class $j$ is given by $\Lambda_{j}=\int_{0}^{1}\lambda_{j}(t)dt$.
@@ 3.2 MDP Model
We use an MDP framework to decide on whether to assign a waiting customer to one of the overflow pools. The objective is to find the optimal overflow policy that minimizes the infinite-horizon, long-run average cost. We consider a discrete-time setup, where the decision maker (e.g., hospital manager) makes overflow decisions at $m$ predefined time epochs each day (e.g., hourly), denoted as $t_{0},t_{1},\dots,t_{m-1}$ for day $t$.
@@
  • $X_{j}(t_{k})$, $j=1,\dots,J$, denotes the customer count for pool $j$. This count includes both the number of customers waiting in buffer $j$ and those in service in pool $j$; we show below on how to recover each part. The former (waiting count) only includes class $j$ customers, while the latter (in-service count) may include customers from other classes due to overflow assignments.
  • $Y_{j}(t_{k})$, $j=1,\dots,J$, denotes the to-depart count in pool $j$. This quantity represents the number of customers anticipated to depart from pool $j$ between the present time and the end of the current day. Tracking this quantity is essential for the two-time-scale inpatient departure scheme, where the “planned discharge” information is used to inform overflow decisions.
@@
To explain, the total number of overflow assignments from all non-primary classes to pool $j$ must not exceed the number of idle servers in pool $j$, which can be computed as $(N_{j}-x_{j})\vee 0$, with $x\vee 0=\max(x,0)$ for any real number $x$. Similarly, the total number of assignments for class $i$, including overflow assignments to a pool $\ell\neq i$ and those kept waiting, must equal the number of waiting customers. This can be calculated as $(x_{i}-N_{i})\vee 0$, which we also refer to as the queue length. We use $q_{i}=(x_{i}-N_{i})\vee 0$ to denote this queue length for class $i$ (buffer $i$). Holding overflow decisions until predefined decision epochs allows managers to use up-to-date system information for better coordination. This practice is common in hospital assignment settings (Bertsimas and Pauphilet 2024, Han et al. 2021) and has also been shown to improve decision quality in other operational contexts by allowing information to accumulate (Xie et al. 2025).
@@
We denote the state at epoch $t_{k}$ before taking an action as $s$, i.e., the pre-action state. After an action is taken, the resulting state is denoted as $s^{+}$, i.e., the post-action state. We further denote by $s^{\prime}$ the state after the exogenous events (random arrivals and departures) occurring between the current epoch $t_{k}$ and the next epoch $t_{k+1}$. The transition from $s$ to $s^{+}$ is the same across all epochs (adding overflow-in patients and subtracting overflow-out patients), whereas the transition from $s^{+}$ to $s^{\prime}$ differs between the midnight epoch ($k=0$) and non-midnight epochs ($k\neq 0$). For non-midnight epochs, given an action $f={f_{i,j}}$, the transition from $s$ to $s^{\prime}$ follows:
$x_{j}^{\prime}$ $=x_{j}^{+}+a_{j}-d_{j}=\left(x_{j}+\sum_{i=1,i\neq j}^{J}f_{i,j}-\sum_{k=1,k\neq j}^{J}f_{j,k}\right)+a_{j}-d_{j};\quad j=1,\dots,J,$ (2)
$y_{j}^{\prime}$ $=y_{j}-d_{j};\quad j=1,\dots,J,$
$h^{\prime}$ $=(h+1)\mod m,$
where $a_{j}$ and $d_{j}$ denote, respectively, the number of class-$j$ arrivals and departures from pool $j$ between the two epochs. The distributions of arrivals and departures are specified in Appendix 9. For the midnight epoch, the main difference in the transition is in the to-depart count, with $y^{\prime}_{j}\sim\text{Bin}(\min\{x_{j},N_{j}\},\mu_{j})$ for $j=1,\dots,J$, meaning that each customer currently in service independently departs on the next day with probability $\mu_{j}$. We denote the transition probability from state $s$ to state $s^{\prime}$, through $s^{+}$ under action $f$, is denoted as $p(s^{\prime}|s,f)$. The detailed calculation of these transition probabilities is provided in Appendix 9.
@@
where $q_{j}^{+}=(x_{j}^{+}-N_{j})\vee 0$ denotes the post-action queue length for class $j$.
@@
We consider the long-run average cost objective since overflow decisions have long-term impacts on the system. We focus on all feasible stationary Markovian policies, and we denote the policy space as $\Pi$. For a policy $\pi\in\Pi$, we use $\pi(f\mid s)$ to denote the probability of selecting a feasible action $f$ in a given state $s$ (where deterministic policies are included via using the Dirac measure). The resulting stochastic process $\{S(t_{k})\}$ forms a discrete-time Markov Chain (DTMC) with a one-epoch transition probability from state $s$ to $s^{\prime}$ as $p_{\pi}(s^{\prime}|s)=\mathbb{E}_{f\sim\pi(\cdot|s)}p(s^{\prime}|s,f)$. The expected one-epoch cost is $g_{\pi}(s)=\mathbb{E}_{f\sim\pi(\cdot|s)}g(s,f)$. Correspondingly, the objective is to identify the optimal policy
@@
While our baseline setting does not consider transferring overflow customers back to their primary pool, this can be accommodated by adding actions that represent the number of in-service transfers from each pool $j$ back to their primary pools. In addition, although we assume a linear cost function in this paper, our framework can handle more complex cost structures, such as polynomial costs or costs related to service levels (e.g., the proportion of patients waiting more than 6 hours). See the appendix of Dai and Shi (2019) for an example to account for service-level metrics. We note that these adjustments do not alter the core elements of our problem, and our algorithmic framework is both flexible and general enough to accommodate them. For ease of exposition, we will focus on the current baseline setting, where primary assignments are handled automatically, departures are pool-dependent, no transfer back of overflow customers, and the cost is linear.
@@ 4 Atomic Actions and Proximal Policy Optimization
One of the core challenges in solving the MDP model developed in Section 3.2 is the combinatorial explosion of the action space. Specifically, when there are enough idle servers in the system, the action space has a size of $\prod_{j=1}^{J}\left(\begin{array}[]{c}q_{j}+w_{j}\\ w_{j}\end{array}\right)$, where $q_{j}$ is the number of waiting customers in class $j$ and $w_{j}$ is the number of overflow pools for class $j$ that have idle servers. To illustrate, in a twenty-pool system with each server-pool having a capacity of 60 and a state $(x_{1},x_{2},x_{3},\dots,x_{20})=(65,63,62,50,50,50,\dots,50)$, the number of potential ways to assign 10 waiting customers can exceed $10^{9}$. Even if the value function is computed (either exactly or via approximation), brute-force searching through all possible actions to find the one with the lowest cost-to-go is infeasible. To overcome this curse of dimensionality, we use atomic actions in combination with PPO that form the key foundation for our algorithm’s scalability. In Section 4.1, we introduce the atomic action and policy setup. In Section 4.2, we present the PPO framework.
@@ 4.1 Action Space Reduction: Atomic Actions
To elaborate, for a given state $s=(x_{1},\dots,x_{J},y_{1},\dots,y_{J},h)$ observed at epoch $t_{k}$, the total number of waiting customers across all classes is $q=\sum_{j=1}^{J}q_{j}$, where $q_{j}=(x_{j}-N_{j})\vee 0$. Following a pre-determined rule (e.g., first-come-first-serve), we determine an order for these $q$ waiting customers and conduct atomic actions $\mathbf{a}=\{a^{1},a^{2},\dots,a^{q}\}$ for each of them sequentially. Here, $a^{n}$ represents the assignment for the $n$th customer in the chosen order. The action space for each $a^{n}$ is simply $\mathcal{J}=\{1,\dots,J\}$. If the class of the $n$th customer is $j$, then $a^{n}=j$ indicates the customer will continue waiting until the next decision epoch; otherwise, the customer is assigned to an overflow pool (see details on feasible overflow assignment in Section 4.1.2).
$^{1}=sss^{\prime}$$\textbf{a}=(a^{1},\dots,a^{q})$$s^{2}$$s^{3}$$s^{q+1}$$\sigma^{1}(s),a^{1}$$\sigma^{2}(s),a^{2}$ArrivalDepartureTime $t_{k}$Time $t_{k+1}$
@@
Figure 1 illustrates the atomic decision process for deciding and executing the atomic actions. We denote the sequence of customer classes of the chosen order as $\boldsymbol{\sigma}(s)=(\sigma^{1}(s),\sigma^{2}(s),\cdots,\sigma^{q}(s))$, with $\sigma^{n}(s)$ representing the class of the $n$th waiting customer within the chosen order. Starting from the first customer, who is from class $\sigma^{1}(s)$, we execute the atomic action $a^{1}$ for this customer, then we update the state according to $a^{1}$ and sequentially move to the second customer, so on and forth till we execute the atomic action $a^{q}$ for the last waiting customer, who is from class $\sigma^{q}(s)$. The atomic action $a^{n}$ for the $n$th customer is chosen based on the atomic state, $s^{n}$, which is observed after assigning the (n1)(n-1)th customer. This state has the same structure as the one introduced in Section 3.2, i.e., $s^{n}=(x^{n}_{1},\dots,x^{n}_{J},y_{1},\dots,y_{J},h)$. After executing $a^{n}$, the atomic decision process updates the state to $s^{n+1}$ as
@@
For clarity, and to distinguish between the MDP state and action defined in Section 3.2 and the ones used for the atomic decision process here, we will refer to those in Section 3.2 as the system-level state and action throughout the rest of the paper. Meanwhile, we will refer to the ones in the atomic decision process as the atomic state and action. It is important to note that the system-level and atomic states and actions are closely related. Specifically, the initial atomic state $s^{1}$ corresponds to the system-level (pre-action) state $s$, and after all atomic actions are executed, the final atomic state $s^{q+1}$ corresponds to the system-level post-action state $s^{+}$. For any given pair $(\boldsymbol{\sigma}(s),\textbf{a})$, we can recover the system-level action $f_{i,j}$ via $\sum_{n=1}^{q}\mathbbm{1}(a^{n}=j,\sigma^{n}(s)=i)$.
Theoretical support for the proposed atomic action approach comes from a contemporary paper Dai et al. (2025b). Their Theorems 1 and 2 show that, in certain stochastic processing networks, (i) decomposing joint actions into atomic steps incurs no loss in optimality, and (ii) even enforcing step-independent policies (each atomic action follows the same rule) still achieves the optimal long-run average reward. This step-independent structure mirrors our setup in Section 4.1, where each atomic action is sampled from the same policy $\pi(\cdot\mid s^{\ell})$ regardless of step index $\ell$.